OpenStudy (anonymous):

Prove by induction: a) n! >= n*2^(n-2) b) n! <= n^(n-1) Thanx!

OpenStudy (anonymous):

okay n! is meant to n(n-1) so we got n(n-1)>= n*2^(n-2) by divide each sides by (n-1) n>=2^(n-2) that's mean is correct subtitde by 2 for example 2>=2^0 2>=1

OpenStudy (anonymous):

but i have to prove it by induction, so (n+1)! >=...

OpenStudy (anonymous):

assume it to be true upto n (n+1)! = n! (n+1) >= (n+1)*2^(n-1) => n! >=2^(n-1) => n * 2^(n-2) for n >= 2 therefore (n+1)! >= (n+1)*2^(n+1-2)

OpenStudy (anonymous):

assume b) upto n ... (n+1)! = n! (n+1) <= (n+1)^n = (n+1) (n+1)^(n-1) => n! <= (n+1)^(n-1) which is true ... therefoe (n)! <= n^(n-1)

OpenStudy (anonymous):

hash the second step how is it thata

OpenStudy (anonymous):

sorry i didn't get u

OpenStudy (anonymous):

n!(n+1) what is equal

OpenStudy (anonymous):

where??

OpenStudy (anonymous):

(n+1)! = (n+1) x n!

OpenStudy (anonymous):

okay and n! can it simplify or not

OpenStudy (anonymous):

b) Base Case: n = 1; 1! = 1 n^(n-1) = 1 ^ 0 = 1 1 <= 1 is true. Base Case holds// Assume n = k holds; hence: k! <= k^(k-1) is true! prove that n = k + 1 holds too. LHS: Now (k+1)! = (k+1)k! RHS: (k+1) ^ k need to prove that: (k+1)k! <= (k+1)^k Divide by k+1 (since its positive): k! <= (k+1)^(k-1) Now since we know that k! <= k^(k-1) is true, the right hand side can only grow with a k+1, hence k! remains <= (k+1)^(k-1)

OpenStudy (anonymous):

aaaah okok, i see now, thnx everybody!

OpenStudy (anonymous):

thanx